OpenStudy (anonymous):

Two triangles are similar and have a ratio of similarity of 8:3. What is the ratio of their perimeters and the ratio of their areas?

OpenStudy (anonymous):

The ratios of perimeters will be 8:3. The ratio of areas will be 64:9. A1 = 1/2bh A2 = (1/2)8/3b*8/3h = (8/3)^2 1/2bh = 64/9 A1